OpenStudy (anonymous):

(x)/b-3 = bx-6 Can someone take me step by step on how to do this? (solve for x)

OpenStudy (the_fizicx99):

You can use b = 2 and x = 2 Or x = -2 and b = -2 2/2 = 1 1-3 = -2 b*x = 2*2 4-6 -2 = -2

OpenStudy (anonymous):

So the answer is 2 ?

OpenStudy (the_fizicx99):

b = 2 x = 2 OR x = -2 and b = -2 I'm assuming you know that 2 negatives cancel out and make a positive xD

OpenStudy (anonymous):

That's not correct

OpenStudy (anonymous):

(x)/b-3 = bx-6 x=(b-3)(bx-6) x=b^2x-6b-3bx+18 x-b^2x+3bx=-6b+18 x(1-b^2+3b)=-6b+18 x=(-6b+18)/(1-b^2+3b)

OpenStudy (anonymous):

0_0

OpenStudy (the_fizicx99):

o.o ohh they're polynomials..? -sigh-